OpenStudy (anonymous):

FIND THE ANSWER IS SLOPE-INTERCEPT AND STANDARD FORM! perpendicular to 3x+4y=13 and through (2,7)

OpenStudy (anonymous):

would 3x+4y=13 be standard form? (Ax+By=C)

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

so then what would i have to plug in to get the slpe-intercept?

OpenStudy (jhannybean):

Slope intercept form: \(\large y=mx+b\)

OpenStudy (jhannybean):

You want to isolate the y on one side and your x's and constants on the other side.

OpenStudy (anonymous):

so: y= -3/4x + 13/4

OpenStudy (jhannybean):

To accomplish this, subtract -3x from both sides and then divide the whole fraction on the RHS by 4. You will have isolated your y.

OpenStudy (jhannybean):

Great job! You got it :)

OpenStudy (anonymous):

thanks!

OpenStudy (jhannybean):

:D smart girl!